OpenStudy (anonymous):

George tells you that when variables are in the denominator, the equation 4/5 + 3/x+1/2 becomes unsolvable. George explains, "There is a value for x that makes the denominator zero, and you can't divide by zero." Demonstrate to George how the equation is still solvable and explain your reasoning.

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@jim_thompson5910

jhonyy9 (jhonyy9):

4/5 + 3/x + 1/2 - so this wan being one equation ? so than you need making equal by zero or by anything - suppose by zero 4/5 +3/x + 1/2 = 0 - so the first restriction for x is x=0 because than the fraction 3/x will be undefined - common denominator will be 10x so than multiplie first numerator by 2x ,second by 10 and 3rd by 5x and will get 8x +30 + 5x = 0 13x +30 =0 13x = -30 x = -30/13 hope this is understandably